# Combining LLM Semantic Reasoning with GNN Structural Modeling for Multi-View Multi-Label Feature Selection

### v1

@@ Introduction
Table 2: Experimental results (mean ± std) in terms of AP and Macro-AUC, where the 1st/2nd best results are shown in boldface/underline.
@@
Table 3: Experimental results (mean ± std) in terms of LRAP and HL, where the 1st/2nd best results are shown in boldface/underline.
@@
Table 4: Ablation experimental results in terms of AP, AUC, LRAP, and HL on SCENE and Yeast.
@@ Conclusion
We propose a novel method for multi-view multi-label feature selection that integrates large language model priors with graph attention mechanisms. By constructing a heterogeneous graph that encodes both LLM-derived semantic information and statistical information based on mutual information as well as label co-occurrence, and applying a lightweight Graph Attention Network, the method produces informative feature embeddings for effective ranking and selection. Extensive experiments on benchmark datasets show consistent improvements over state-of-the-art methods, demonstrating superior performance, robustness, and generalization even on small-scale datasets. In future work, we will explore end-to-end optimization with LLM feedback for closed-loop refinement, integration of learned embeddings into downstream tasks, and self-supervised pretraining on the heterogeneous graph to enhance representation quality under limited supervision.

### v2

@@ Introduction
Table 2: Experimental results (mean ± std) for AP and AUC, where the 1st/2nd best results are shown in boldface/underline.
@@
Table 3: Experimental results (mean ± std) for LRAP and HL, where the 1st/2nd best results are shown in boldface/underline.
@@
Table 4: Ablation experimental results on SCENE and Yeast.
@@ Conclusion
We propose a novel method for multi-view multi-label feature selection that integrates large language model priors with graph attention mechanisms. By constructing a heterogeneous graph that encodes both LLM-derived semantic information and statistical information based on mutual information as well as label co-occurrence, and applying a lightweight Graph Attention Network, the method produces informative feature embeddings for effective ranking and selection. Extensive experiments on benchmark datasets show consistent improvements over state-of-the-art methods, demonstrating superior performance, robustness, and generalization even on small-scale datasets. In future work, we will explore end-to-end optimization with LLM feedback for closed-loop refinement, integration of learned embeddings into downstream tasks, and self-supervised pretraining on heterogeneous graphs to enhance representation quality.